# Neural Spacetimes for DAG Representation Learning

### v1

@@ 5 Conclusion
Acknowledgements
AK acknowledges financial support from an NSERC Discovery Grant No. RGPIN-2023-04482 and No. DGECR-2023-00230. AK also acknowledges that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute [^3] https://vectorinstitute.ai/partnerships/current-partners/.
Appendix A Additional Background
A.1 Dimension and Size of a Metric Spaces
Definition 5 (Doubling Constant).
The doubling constant $K\geq 1$ of a metric space $(X,d)$ is the smallest integer $k\geq 1$ for which the following holds: for every center $x\in X$ and each radius $r>0$, every (open) ball $B(x,r)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{u\in X:d(x,u)<r\}$ can be covered by at most $k$ (open) balls of radius $r/2$; i.e., there exist $x_{1},\dots,x_{k}\in X$ such that $B(x,r)\subseteq\bigcup_{i=1}^{k}B(x_{i},r/2)$.
Definition 6 (Diameter).
Let $(X,d)$ be a metric space. The diameter of $(X,d)$ is defined to be $\operatorname{diam}(X,d)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{sup}_{x,\tilde{x}\in X}\,d(x,\tilde{x})$.
Definition 7 (Separation).
Let $(X,d)$ be a metric space. If $X$ has at-least two points, then its separation is defined to be $\operatorname{sep}(X,d)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{inf}_{x,\tilde{x}\in X,\,x\neq\tilde{x}}\,d(x,\tilde{x})$, otherwise $\operatorname{sep}(X,d)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1$.
A.2 Invertible Positive Matrices
Definition 8 (Invertible Positive Matrices $I^{+}_{D}$).
$\mathsf{W}=\lambda\,I_{D}+|\tilde{\mathsf{W}}|,$
A.3 Neural Snowflakes
Neural Snowflakes. Neural snowflakes leverage quasi-metric spaces and implement a learnable adaptive geometry $\|\hat{x}_{u}-\hat{x}_{v}\|_{f}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}f(\|\hat{x}_{u}-\hat{x}_{v}\|),$ where $\hat{x}_{u},\hat{x}_{v}\in\mathbb{R}^{D}$. In particular, it is a map $f:[0,\infty)\rightarrow[0,\infty)$, with iterative representation
$f(t)$ $=u_{J}^{1+|p|}$ (10)
[…]
$u_{0}$ $=u$
$\psi_{a,b}(u)_{k}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{(}1-e^{-|u_{k}|},|u_{k}|^{a},\log(1+|u_{k}|)^{b}\big{)},$ (11)
A.4 Pseudo-Riemannian Manifolds and Lorentzian Spacetimes
Pseudo-Riemannian manifolds are a generalization of Riemannian manifolds where the metric tensor $g$ is not constrained to be positive definite. A $d$-dimensional pseudo-Riemannian manifold is a smooth manifold $M$ equipped with a pseudo-Riemannian metric $g$. Formally, a pseudo-Riemannian metric tensor is a smooth, symmetric, non-degenerate bilinear form (called a scalar product) defined on the tangent bundle $TM$ (disjoint union of all tangent spaces). The metric tensor a each point $p\in M$ is denoted by $g_{p}$. The metric assigns to each tangent space $T_{p}M$ a signature (±,±,,±)plus-or-minusplus-or-minusplus-or-minus($\pm$,$\pm$,$\dots$,$\pm$)( ± , ± , … , ± ). In particular, the tangent space $T_{p}M$ admits an orthonormal basis $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{d}\}$ that satisfies $\forall i\in\{1,\dots,d\},~{}g_{p}(\mathbf{e}_{i},\mathbf{e}_{i})=\pm 1$. If $g_{p}(\mathbf{e}_{i},\mathbf{e}_{i})=1$ for all $i$, then $M$ is a Riemannian manifold, its metric tensor is positive definite, and its metric signature is (+,+,,+)(+,+$\dots$,+)( + , + , … , + ).
A.5 Related Work on Spacetime Representation Learning
A.6 Hyperbolic Spaces as Compared to Neural Spacetime Embeddings
Appendix B Proofs
B.1 Properties of Neural Spacetimes
Proof of Proposition 1.
Fix $T\in\mathbb{N}_{+}$ and set $D=0$. Fix a $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ with representation equation 4. We verify that the relation $\lesssim^{\mathcal{T}}$ is a partial order on $\mathbb{R}^{D+T}$. It is trivially symmetric since $T(x)_{t}=T(x)_{t}$ for each $t=1,\dots,T$ and every $x\in\mathbb{R}^{D+T}$; thus, $x\lesssim^{\mathcal{T}}x$ for each $x\in\mathbb{R}^{D+T}$.
Since each matrix $V_{j}\in I_{T}^{+}$ then, as shown in equation 12, it is injective. Since the leaky ReLU function is injective, and so is its componentwise composition, as well as affine translations/shifts, then $\mathcal{T}$ is injective. Thus, for each $y,x\in\mathbb{R}^{D+T}$ if $\mathcal{T}(x_{t})_{t}=\mathcal{T}(y_{t})_{t}$ for each $t=1,\dots,T$ then $\mathcal{T}(x)=\mathcal{T}(y)$; by injectivity of $\mathcal{T}$ [^4] Since $D=0$. we conclude that $x=y$.
The transitivity follows directly from the componentwise $\leq$ (i.e. the product order on $\mathbb{R}^{T}$). ∎
Proof of Proposition 2.
Symmetry of $\mathcal{D}$:
The map $\mathcal{D}$ is symmetric since
Non-Negativity of $\mathcal{D}$:
$(\mathsf{W}u)_{i}=\sum_{j=1}^{D}\,\mathsf{W}_{i,j}u_{j}\geq 0.$
Thus, $\mathcal{D}(x,y)\geq 0$ for each $x,y\in\mathbb{R}^{D}$.
Case I : $\mathcal{D}$ Separates Points if $T=0$:
$\mathcal{D}(x,y)$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathsf{W}_{J}\sigma_{s_{J+1},l_{J+1}}(u_{J})$
$u_{j}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathsf{W}_{j}\sigma_{s_{j},l_{j}}\bullet(u_{j-1})\qquad\qquad\mbox{ for }j=1,\dots,{J-1}$
$u_{0}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}|\sigma_{s_{0},l_{0}}(x)-\sigma_{s_{0},l_{0}}(y)|$
$\sigma_{s,l}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}\operatorname{sgn}(x)\,|x|^{s}&\mbox{ if }|x|<1\\ \operatorname{sgn}(x)\,|x|^{l}&\mbox{ if }|x|\geq 1\end{cases}.$
$(\mathsf{W}_{1})_{i,i}=\lambda+\sum_{j=1}^{D}\,|\tilde{\mathsf{W}}_{i,j}|>\sum_{j=1}^{D}\,|\tilde{\mathsf{W}}_{i,j}|\geq\sum_{j=1;\,j\neq i}^{D}\,|\tilde{\mathsf{W}}_{i,j}|\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}R_{i},$ (12)
$\Lambda_{1}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcup_{i=1}^{D}\,\overline{\operatorname{B}}_{2}\big{(}(\mathsf{W}_{1})_{i,i},R_{i}\big{)}$
where, for $u\in\mathbb{C}$ and $r\geq 0$ we define $\overline{\operatorname{B}}_{2}\big{(}u,r\big{)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{z\ \in\mathbb{C}:\,\|u-z\|\leq r\}$. Since the computation in equation 12 showed that $R_{i}<(\mathsf{W}_{1})_{i,i}$ for each $i=1,\dots,D$ then $0\not\in\Lambda_{1}$. Thus, $\mathsf{W}_{1}$ is invertible.
Case II: Pseudo-Metric for Positive Time Dimensions if $T>0$:
Relaxed Triangle Inequality:
$\beta_{j}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\min\{(\max\{s_{j},t_{j}\}-1),0\}.$
Note that, $\beta_{j}=0$ whenever both $s_{j}$ and $t_{j}\leq 1$ and it equals to $(\max\{s_{j},t_{j}\}-1)$ otherwise. Further, note that if $s_{j}=t_{j}>1$, then $\beta_{j}=s_{j}-1$.
$\|u_{j}\|\leq\|\mathsf{W}_{j}\|_{op}\,\|\sigma_{s_{j},l_{j}}(u_{j-1})\|.$ (13)
By (Xia, 2009, Example 2.2): for each $j=1,\dots,J$ we have that
$\|\sigma_{s_{j},l_{j}}(u_{j-1})\|\leq 2^{\beta_{j}}\,\|u_{j-1}\|.$ (14)
[…]
Again using (Xia, 2009, Example 2.2) we have that
$=$ $2^{\beta_{0}}\,\Biggl{(}\sum_{i=1}^{d}\,|x_{i}-y_{i}|^{2}\Biggr{)}^{1/2}$
$=$ $2^{\beta_{0}}\,\|x-y\|.$
[…]
Finally, notice that if for each $j=1,\dots,J$ the matrix $\mathsf{W}_{j}$ is orthogonal and $0<s_{j},t_{j}\leq 1$ then equation 18 becomes $1\,\|x-y\|$; in which case $\mathcal{D}$ is a metric. This concludes the proof. ∎
B.2 Embedding Results
We will routinely use the following ordering.
Definition 9 (Product Order).
Equivalently, $x\lesssim^{\times}y$ if $1=\prod_{t=1}^{T}\,I_{x_{t}\leq y_{t}}$.
Lemma 1 (An Implementation Lemma for Neural Spacetimes).
  • (i)
    Implementation of Product Ordering:
    $(x)_{D+1:D+T}\lesssim^{\times}(y)_{D+1:D+T}\Leftrightarrow\mathcal{T}(x)_{t}\leq\mathcal{T}(y)_{t}\,\,\mbox{ for }\,t=D+1,\dots,D+T$
  • (ii)
    Implementation of Snowflake of the $\ell^{p}$ metric:
    $\mathcal{D}(x,y)=\|(x)_{0:D}-(y)_{0:D}\|_{p}^{\alpha}$
  1. (a)
  2. (b)
  3. (c)
    No. Non-zero Parameters: $\operatorname{No.Param}(\mathcal{T})=T+2$ and $\operatorname{No.Par}(\mathcal{D})=3(2+D)$.
Note that in the main text, we use $\lesssim^{\mathcal{T}}$ to indicate more explicitly that the order $\lesssim^{\times}$ is implemented according to the neural partial order $\mathcal{T}$.
Proof of Lemma 1.
Step 1 - Implementation of the Product Order
$\mathcal{T}(x)$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}I_{T}\sigma_{1,1}\bullet(x)_{D+1:D+T}+\mathbf{0}_{T}$ (19)
Fix $J=2$, set $s_{0}=l_{0}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1$, $s_{1}=l_{1}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}p$, $\mathsf{W}_{1}=I_{D}$, and $s_{2}=l_{2}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1/(p\alpha)$, $\mathbf{1}_{D}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathsf{W}_{2}\in\mathbb{R}^{1\times D}$ with $(\mathsf{W}_{2})_{i}=1$ for each $i=1,\dots,D$. Therefore,
$u_{0}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{(}|\sigma_{1,1}(x_{i})-\sigma_{1,1}(y_{i})|\big{)}_{i=1}^{d}=\big{(}|x_{i}-y_{i}|\big{)}_{i=1}^{d}$ (20)
$\therefore\,\,u_{1}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}I_{D}\sigma_{p,p}\bullet(u_{0})=\big{(}|x_{i}-y_{i}|^{p}\big{)}_{i=1}^{d}$
$\therefore\,\,u_{2}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma_{\alpha/p,\alpha/p}\big{(}\mathbf{1}_{D}u_{1}\big{)}=\biggl{(}\sum_{i=1}^{D}\,|x_{i}-y_{i}|^{p}\bigg{)}^{\frac{\alpha}{p}}=\|(x)_{1:D}-(y)_{1:D}\|_{p}^{\alpha}.$
Consequentially, the map $\mathcal{D}(x,y)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}u_{2}$, with $u_{2}$ defined by equation 20, is of the form of equation 3 and satisfies: for each $x,y\in\mathbb{R}^{D+T}$
$\mathcal{D}(x,y)=\|(x)_{1:D}-(y)_{1:D}\|_{p}^{\alpha}.$
Observe that $\mathcal{D}$ has depth $2$, width $d$, and $2\cdot 3+D+D=2(3+D)$ (non-zero) trainable parameters. ∎
Theorem 3 (Universal Spacetime Embeddings).
Fix $W,N,k\in\mathbb{N}_{+}$, $K>0$, and let $(P,\lesssim,d)$ be a $k$-point causal metric space with doubling constant $K$, width $W$, and feature encoding $P\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x_{v}\}_{v\in V}\subseteq\mathbb{R}^{N}$. There exists a $D\in\mathcal{O}(\log(K))$, $T\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}W$, an MLP $\mathcal{E}:\mathbb{R}^{N}\to\mathbb{R}^{D+T}$, $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ with representation equation 4, and $\mathcal{D}:\mathbb{R}^{D+T}\times\mathbb{R}^{D+T}\to[0,\infty)$ with representation equation 3 such that: for each $x_{u},x_{v}\in P$
  1. (i)
  2. (ii)
Furthermore, there is some $D\leq k$ such that (ii) can be improved to
$d(x_{u},x_{v})=\mathcal{D}\big{(}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\big{)}$
  1. (i)
    $D+T+8$
    non-zero parameters,
  2. (ii)
    Encoding Complexity: $\mathcal{E}$ depends on
    non-zero parameters.
non-zero parameters.
Proof of Theorem 1.
Step 1 - Causal Embedding:
By (Dilworth, 1950, Theorem 1.1), $(P,\precsim)$ has width $W$ only if there exists an order embedding ~$\tilde{\mathcal{E}}^{\star}:P\to(\{0,1\}^{W},\lesssim^{\times})$. Since the inclusion $\iota$ of ($(\{0,1\}^{W},\lesssim^{\times})$ into $(\mathbb{R}^{W},\lesssim^{\times})$ trivially defines an order embedding, then $\mathcal{E}^{(1)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\iota\circ\tilde{\mathcal{E}}:(P,\precsim)\to(\mathbb{R}^{W},\lesssim^{\times})$ is an order embedding.
Step 2 (Case I) - Metric Embedding - Low-Distortion Case:
where [^5] Note that $\varepsilon\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1-\alpha$ in the notation of Naor & Neiman (2012). $D\in\mathcal{O}(\log(K)/\delta)$. Equivalently, for each $x_{u},x_{v}\in P$
Step 2 (Case II) - Metric Embedding - Isometric Case:
Since $(P,d)$ is such that $P$ is finite, i.e. $k=\#P<\infty$ then,  [^6] The isometric embeddability into a Euclidean snowflake characterizes finite metric spaces; see (Le Donne et al., 2018, Corollary 2.2). implies that setting $\alpha\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\gamma(k-1)/2\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log_{2}(1+1/(k-1))/2\in(0,1]$ (with the case where $\alpha=1$ only being achieved when $P$ is a singleton) there exists some $\tilde{D}\in\mathbb{N}_{+}$ and ~$\tilde{\mathcal{E}}^{(2)}:P\to\mathbb{R}^{\tilde{D}}$ such that: for each $x_{u},x_{v}\in P$
Since $D\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{dim}\big{(}\operatorname{span}\{\tilde{\mathcal{E}}(x_{u})\}_{x_{u}\in P}\big{)}\leq\#P=k$ and since all $D$ dimensional linear of $\mathbb{R}^{\tilde{D}}$ subspaces are isometrically isomorphic to the $D$-dimensional Euclidean space; then there exists some linear surjection $T:\mathbb{R}^{\tilde{D}}\to\mathbb{R}^{D}$ which restricts to a bijective isomorphism from $\operatorname{span}\{\tilde{\mathcal{E}}(x_{u})\}_{x_{u}\in P}$ to $\mathbb{R}^{D}$ (both equipped with Euclidean metrics). Then, the composite map $\mathcal{E}^{(2)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}T\circ\tilde{\mathcal{E}}^{(2)}:(P,d^{\alpha})\to(\mathbb{R}^{D},\|\cdot\|)$ is an isometric embedding. Consequentially: for each $x_{u},x_{v}\in P$ we have that
and we emphasize that $D\leq k$.
Step 3 - Interpolation:
$\mathcal{E}(x_{u})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{(}\mathcal{E}^{(1)}(x_{u}),\mathcal{E}^{(2)}(x_{u})\big{)}.$ (25)
^$\hat{\mathcal{E}}(x_{u})=\mathcal{E}(x_{u}).$ (26)
$\operatorname{aspect}(P,d)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{\max_{x_{u},x_{v}\in P}\,d(x_{u},x_{v})}{\min_{x_{u},x_{v}\in P\,;\,x_{u}\neq x_{v}}\,d(x_{u},x_{v})}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{\operatorname{diam}(P,d)}{\operatorname{sep}(P,d)}.$
Retroactively set $\delta\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1/4\in(0,1]$ and $\alpha\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}3/4\in(1/2,1)$. Then, equation 27 and equation 28 become
In case II, instead,  equation 24 implies that: for each $u,v\in V$
Proof of Proposition 3.
Set $W\in\{1,2\}$ and $D\in\mathbb{N}_{+}$. By definition, of the order on $\mathbb{R}^{D+W}$, given in Lemma 1 (i), a spacetime embedding $\mathcal{E}:P\to\mathbb{R}^{D+W}$ exists if and only if $p\circ\mathcal{E}:P\to\mathbb{R}^{W}$ is an order embedding into $(\mathbb{R}^{W},\lesssim^{\times})$ where $\lesssim^{\times}$ is the product order and $p:\mathbb{R}^{D+W}\ni(x_{i})_{i=1}^{D+W}\to(x_{i})_{i=1}^{W}\in\mathbb{R}^{W}$ is the canonical projection. By (Baker et al., 1970, Theorem 6.1) an order embedding of $(P,\precsim)$ into $(\mathbb{R}^{2},\lesssim^{\times})$ exists if and only if $(P,\precsim)$ has a planar Hasse diagram. ∎
Proof of Theorem 2.
Step 1 - Time Embedding
$\mathcal{E}^{(1)}:(P,\lesssim)\to(\mathbb{R}^{2},\lesssim^{\times}).$
Step 2 - Spatial Embedding
Furthermore, by remark at the beginning of (Rao, 1999, Section 4) by the Johnson-Lindenstrauss lemma, as formulated in (Dubhashi & Panconesi, 2009, Theorem 2.1), one may take $k\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\#P$ an incur an additional factor of $\tilde{c}\sqrt{\log(k)}$, for an absolute constant $\tilde{c}>0$, in the distortion in equation 34; that is find that: for each $u,v\in P$
where $C\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}c\tilde{c}>0$. Set $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1$ and recall that $\|\cdot\|_{2}\leq\|\cdot\|_{1}\leq\tilde{C}\sqrt{\log(k)}\|\cdot\|_{2}$ on $\mathbb{R}^{\tilde{C}\log(k)}$, for any $\tilde{C}>0$. Thus, equation 35 implies that: for each $x_{u},x_{v}\in P$ we have
where $C^{\prime}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}C\tilde{C}>0$ and $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1$.
Step 3 - Interpolation Embedding
Pick some $x^{\star}\in\mathbb{R}^{2+d}\setminus[\cup_{i=1}^{2}\,\mathcal{E}^{(i)}(P)].$ Since $P\subset\mathbb{R}^{N}$ then consider the map $\mathcal{E}:\mathbb{R}^{N}\to\mathbb{R}^{2+d}$ defined for each $x\in\mathbb{R}^{N}$ by
$\mathcal{E}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}\big{(}\mathcal{E}^{(1)}(x),\mathcal{E}^{(2)}(x)\big{)}&\,\mbox{ if }x\in P\\ x^{\star}&\mbox{ if }x\not\in P.\end{cases}$
^$\hat{\mathcal{E}}(x_{u})=\mathcal{E}(x_{u}).$ (37)
$\operatorname{aspect}(P,d_{\operatorname{H}})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{\max_{x_{u},x_{v}\in P}\,d_{\operatorname{H}}(x_{u},x_{v})}{\min_{x_{u},x_{v}\in P\,;\,x_{u}\neq x_{v}}\,d_{\operatorname{H}}(x_{u},x_{v})}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{\operatorname{diam}(P,d_{\operatorname{H}})}{\operatorname{sep}(P,1)}=\operatorname{diam}(P,d_{\operatorname{H}}),$
Since $\mathbb{R}^{2+D}$ is equipped with the product order on its last two (time) dimensions and the $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1/\alpha$-snowflake of the Euclidean $(\ell^{2})$ metric on its first $D$ (space) dimension, then the conditions for Lemma 1 are met. Applying Lemma 1 concludes the proof. ∎
Appendix C Computational Implementation of Neural Spacetimes
C.1 What makes an Activation Function Fractalesque?
The terminology snowflake arises as follows. Consider the activation function $\sigma:\mathbb{R}\to\mathbb{R}$ given by $\sigma(x)=|x|^{\alpha}$ and set $\alpha=\log(4)/\log(3)$. Then, $d_{\alpha}(x,y)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma(|x-y|)$ defines a metric on $\mathbb{R}$ and the metric space $(\mathbb{R},d_{\alpha})$ is isometric to the (von) Koch snowflake fractal (Figure 4) $\mathcal{X}\subset\mathbb{R}^{2}$ endowed with the distance obtained by restriction of the Euclidean distance on $\mathbb{R}^{2}$ to $\mathcal{X}$. Here, $\alpha$ is chosen such that $(\mathcal{X},d_{\alpha})$ has positive and finite $\alpha$-Hausdorff measure.
Indeed, for any general $\alpha\in(0,1)$ one can show that $d_{\alpha}$ defines a metric on $\mathbb{R}$ with the property that any line segment has infinite length, see Semmes (1996). Thus, intuitively, the snowflake space $(\mathbb{R},d_{\alpha}$ contains infinitely more space to place points in while also maintaining a similar geometry to the Euclidean line; see e.g. (Acciaio et al., 2024, Lemma 7.1). See Tyson & Wu (2005) for an intrinsic characterization of metric spaces which arise as snowflakes of a subset of a (possibly multidimensional) Euclidean space.
$|\sigma(x)-\sigma(y)|\leq\omega(|x-y|).$ (38)
$|\sigma(x)-\sigma(y)|\leq\omega(|x-y|)\leq L\,|x-y|^{\alpha}.$ (39)
C.1.1 The activation function in equation 2
Neural spacetimes leverage the following equation:
$\sigma_{s,l}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}\operatorname{sgn}(x)\,|x|^{s}&\mbox{ if }|x|<1\\ \operatorname{sgn}(x)\,|x|^{l}&\mbox{ if }|x|\geq 1,\end{cases}$
(b) Neural (quasi-)metric activation function derivative.
(d) Activation function used by neural partial order (equation 4).
Figure 5: NST activation visualizations.
C.2 Algorithmic Description
[…]

### v2

@@ D.4 Additional Results
D.5 Further Clarifications on Experiments Using Neural Spacetime Embeddings Compared to Fixed-Geometry Baselines
In this section, we aim to provide additional intuition to the reader on how NSTs compare to the fixed-geometry embedding baselines and why we expect them to perform better.
For all experiments, we use a feature encoder that maps the graph node features to the relevant manifold. We assume that the output of the feature encoder is in Euclidean space in all cases. We then proceed according to the specific geometry. This approach aligns with previous works such as neural snowflakes (Borde & Kratsios, 2023) and other studies on product manifold embeddings (Borde et al., 2023b).
Intuitively, one can understand the problem as a two-step process, which in practice is learned jointly via end-to-end gradient descent. In the first step, the encoder learns to position the graph nodes in space—that is, it learns to map the original node features to coordinates in the manifold. In the second step, the distance between points on the manifold is evaluated, either based on a given metric when the geometry is fixed or by learning the geometry (the metric itself) in a data-driven fashion.
In the cases of Minkowski and De Sitter space, the geometry is not learned but given. More specifically, in Minkowski space, we use the feature outputs of the Euclidean feature encoder, apply a change in the metric tensor, and recognize that one coordinate has a different signature. In this particular case, we do not require an exponential map since the space is flat. For De Sitter space, the situation is similar: only the positioning of points in the manifold is optimized via gradient descent, while the geometry remains fixed. In terms of implementation and metric calculation, De Sitter space is a curved manifold, making computations more complicated than for Minkowski space. We use an embedding approach that avoids exponential maps: we map points into a De Sitter hyperboloid in a higher-dimensional Minkowski space and compute geodesic distances there. The geometric operations we utilize include normalizing spatial components, computing the De Sitter inner product, and handling both timelike and spacelike separations. Additionally, we would like to highlight that, although the authors in (Law & Lucas, 2023) also use Minkowski and De Sitter spaces, our baselines are more powerful. This is because we employ a neural network feature encoder to optimize the placement of node embeddings within the manifold
NSTs are intrinsically different in that their geometry is not fixed but rather random at initialization and parametrized by a neural network that always outputs a pseudo-quasi-metric by construction. The specific pseudo-quasi-metric it generates is learned during optimization based on the input graph. Hence, in this case, both the position of the points on the manifold and the geometry of the manifold itself are learned jointly via gradient descent. For NSTs, we do not need exponential maps either since the construction is based on warping a norm. During optimization, the inputs are the graph node features, and the targets are both the distances between nodes (induced by the graph geometry) and the causality structure (given by the edge directions). Since neural spacetimes are able to reshape the geometry of the embedding space as a function of the data, they are inherently more flexible and they are able to embed DAGs with less distortion (the best possible value for distortion is 1).
D.6 Downstream Application Testing: Node classificaition on Embedded Heterophilic graphs
Lastly, we evaluate the downstream performance of our encoding approach in transductive node classification for heterophilic graphs, specifically using the Cornell, Texas, and Wisconsin datasets. It is important to note that this is not a standard application of NSTs, as their original design aims to embed DAGs with minimal distortion. NSTs typically use the feature encoder network as an initial intermediate mapping to a latent space, which is then utilized by the learnable quasi-metric and learnable partial order to encode distance and directionality respectively. While the neural quasi-metric takes two node feature vectors as input and outputs a scalar distance value (making it unsuitable as direct input for a downstream node classification task), the neural partial order operates pointwise and returns a feature vector. Therefore, we adopt the following approach. First, we train the complete NST model to encode the original graph into the continuous geometry, simultaneously optimizing all three networks composing the NST, $\mathcal{E}$, $\mathcal{D}$, and $\mathcal{T}$. Once trained, we use the frozen feature encoder and neural partial order as pretrained featurizers to encode the original node features from the dataset at hand. These transformed features are then fed into a downstream network that is trained as a node-wise classifier. That is, the new node features for the graphs become:
$z_{v}=x_{v}\parallel(\hat{x}_{v})_{1:D}\parallel t_{v},\quad\text{where}\quad\hat{x}_{v}=\mathcal{E}(x_{v}),\;t_{v}=\mathcal{T}(\hat{x}_{v}),$ (43)
where we have augmented the original features with NST-based features optimized according to the graph topology. In the equation above $x_{v}$ corresponds to the original features for node $v$, $\mathcal{E}$ and $\mathcal{T}$ are the frozen feature encoder and partial order, which were originally optimized alongside $\mathcal{D}$. As in the main text, the subscript in $(\hat{x}_{v})_{1:D}$ means that we utilize the first $D$ spatial dimensions from the feature vector, and we use ||||| | for concatenation. We use pre-trained checkpoints with $D=T=10$.
We display the downstream results in Table 11 below. We experiment with different downstream classifiers, including MLPs, GCNs (Kipf & Welling, 2017), GATs (Veličković et al., 2018), CPGNN-MLPs (Zhu et al., 2021), and CPGNN-Cheby (Zhu et al., 2021; Defferrard et al., 2016). Note that CPGNN is not just a network but also a framework for training on heterophilic graphs. It employs an estimator network, such as an MLP or a Chebyshev polynomial-based model, for prediction. The CPGNN method incorporates estimator pretraining and utilizes a compatibility matrix, which is initialized using the training node labels and the training mask. This matrix is further refined using the Sinkhorn-Knopp algorithm for initialization. Furthermore, the compatibility matrix is regularized during training through an additional loss term on top of the cross-entropy loss. We reimplement all baselines and test them on the Cornell, Texas, and Wisconsin datasets using 10-fold splits, based on the masks provided in PyTorch Geometric. Following (Zhu et al., 2021) we use weight decay of $5\times 10^{-4}$ for all networks and pretrain the estimators in CPGNN-MLP and CPGNN-Cheby for 100 steps. All our downstream classifiers use two layers with a ReLU activation function and are trained with a learning rate of 0.01 for 400 steps using AdamW. We experiment with varying hidden dimensions—10, 20, 30, and 64—the last of which is the hidden dimension used by the baselines in (Zhu et al., 2021). In our experiments, we study the effect of adding geometric node features based on the NST encoding to downstream classifiers. We find that the MLP classifier achieves the best downstream performance when augmented with NST features, particularly for the Texas and Wisconsin datasets. For Cornell, the best downstream performance is achieved by the MLP without NST features, which attains $71\%$ accuracy with a hidden dimension of 64. The counterpart augmented with NST features achieves $70\%$ accuracy, making the difference not particularly significant. Overall, in Cornell, we observe that adding NST features is especially beneficial for smaller classifiers with fewer hidden dimensions.
These experiments confirm that the NST has learned meaningful features that can be used by downstream classifiers. However, we invite future research in this direction, as this paper primarily focuses on embeddings, and this appendix is only an initial exploration of the applicability of NSTs to downstream tasks.
Table 11: Node classification results for heterophilic graphs, including the test accuracy mean and standard deviation. ✓ denotes that the input features to the downstream classifier consist of the original graph input features plus the NST features $z_{v}=x_{v}\parallel(\hat{x}_{v})_{1:D}\parallel t_{v}$. In contrast, ✗ indicates that only $x_{v}$ is passed into the model.
Dataset Hidden Dim NST Features MLP GCN GAT CPGNN-MLP CPGNN-ChebNet
Cornell 10 0.55 $\pm$ 0.09 0.43 $\pm$ 0.09 0.48 $\pm$ 0.08 0.54 $\pm$ 0.05 0.51 $\pm$ 0.05
10 0.40 $\pm$ 0.12 0.43 $\pm$ 0.06 0.48 $\pm$ 0.07 0.47 $\pm$ 0.10 0.46 $\pm$ 0.11
20 0.61 $\pm$ 0.09 0.44 $\pm$ 0.09 0.49 $\pm$ 0.08 0.51 $\pm$ 0.05 0.50 $\pm$ 0.10
20 0.56 $\pm$ 0.11 0.43 $\pm$ 0.06 0.49 $\pm$ 0.09 0.43 $\pm$ 0.11 0.41 $\pm$ 0.07
30 0.70 $\pm$ 0.06 0.44 $\pm$ 0.09 0.47 $\pm$ 0.09 0.47 $\pm$ 0.12 0.52 $\pm$ 0.07
30 0.70 $\pm$ 0.05 0.43 $\pm$ 0.07 0.50 $\pm$ 0.10 0.42 $\pm$ 0.11 0.41 $\pm$ 0.11
64 0.70 $\pm$ 0.07 0.43 $\pm$ 0.09 0.50 $\pm$ 0.06 0.50 $\pm$ 0.08 0.52 $\pm$ 0.07
64 $\mathbf{0.71\pm 0.07}$ 0.43 $\pm$ 0.07 0.50 $\pm$ 0.08 0.40 $\pm$ 0.10 0.40 $\pm$ 0.11
Texas 10 0.58 $\pm$ 0.14 0.58 $\pm$ 0.08 0.56 $\pm$ 0.09 0.61 $\pm$ 0.05 0.62 $\pm$ 0.08
[…]